OpenStudy (anonymous):

Using trigonometric identities, transform the left side of the equation into the right side: tan(x)cot(x) = 1

OpenStudy (***[isuru]***):

Hi, cot(x) is the reciprocal of tan(x) which means, cot(x) = 1/tan(x) So let's keep it in our mind and move to the question tan(x) * cot(x) = 1 L.H.S = tan(x) * cot(x) [ let's substitute cot(x) = 1/tan(x) for cot(x) ; ] = tan(x) * 1/tan(x) = tan(x)/tan(x) = 1 = R.H.S That's it!